Solve the following word problem (Hint: setup and solve a system of equations):
One number plus twice another number gives a sum of 42. The difference of these same two numbers equals –3.
Find the two numbers

Answer by mananth(16946)

let one number be x & the other be y
x+2y= 42
x-y= -3
..
(x+2y)-(x-y)=42-(-3)
x+2y-x+y=45
3y=45
y=9
x=6